Assume that the six month treasury spot rate is 1.6% APR and the one-year rate is 2% APR, both compounded semiannually. What is the price of a one year $1,000 par treasures bond with 2% coupons?
Answer #1

Price of the bond is the present value of future payments.

Price = (Coupon/(1+rate)) + (Future value /(1+rate))

six months rate = 1.6%/2 = 0.8%

1 year rate = (1+(2%/2))^2 -1 = 2.01%

Semiannual coupon = 1000*2%/2 =$10

-------------------------------------------------------------

Price = (10/(1+0.8%)) + ((1000+10)/(1+2.1%))

Price = 9.9206 + 990.099

Price = 1000.02